Margaritaville Hotel Properties is opening a new beach resort in Tybee Island, GA at a cost of $250 Million in year 0. The hotel is expected to operate for 20 years and at the end, be sold for approximately $500 Million in year 20. As an investment the hotel expected to earn $27 Million per year (including $20 Million in year 20). With a discount rate of 8% and reinvestment rate of 8%, analyze the projects feasibility using:

Payback 

Discounted Payback 

NPV 

IRR 

Profitability Index 

MIRR

It turns out, you forgot that the franchise company that licenses the Margaritaville name will require the hotel owners to renovate the hotel property in year 10. This will result in significant room closures and a significant capital investment. Therefore, cash flows in that year 10 are expected to be -$5 Million. Using MIRR, what did you get for a rate of return?
Answer #1

The cash flows and cumulative CF are as follows

Year Initial cost Annual earnings Salvage Net Cash flow Cumulative CF Discounted CF Cumulative CF
0 -250 -250 -250 -250.00 -250.00
1 27 27 -223 25.00 -225.00
2 27 27 -196 23.15 -201.85
3 27 27 -169 21.43 -180.42
4 27 27 -142 19.85 -160.57
5 27 27 -115 18.38 -142.20
6 27 27 -88 17.01 -125.18
7 27 27 -61 15.75 -109.43
8 27 27 -34 14.59 -94.84
9 27 27 -7 13.51 -81.33
10 27 27 20 12.51 -68.83
11 27 27 47 11.58 -57.25
12 27 27 74 10.72 -46.53
13 27 27 101 9.93 -36.60
14 27 27 128 9.19 -27.41
15 27 27 155 8.51 -18.89
16 27 27 182 7.88 -11.01
17 27 27 209 7.30 -3.72
18 27 27 236 6.76 3.04
19 27 27 263 6.26 9.30
20 20 500 520 783 111.57 120.86

As per rules I am answering the first 4 subparts of the question

Payback = 9+ 7/27 = 9.26 years

Discounted payback = 17 + 3.72/6.76 = 17.55 years

NPV = $120.86 m

IRR= 12.13%

WORKINGS